OpenStudy (allie2020):

@jtug3 Which formula gives the zeros of y = sin(x)? k pi for any positive integer k k pi for any integer k k pi /2 for any positive integer K pi/ 2 for any integer k

OpenStudy (allie2020):

will award medal.

OpenStudy (allie2020):

@YanaSidlinskiy have any idea?

OpenStudy (mayankdevnani):

\[\large \bf \sin(- \pi)=-\sin \pi=0\]

OpenStudy (mayankdevnani):

well option B is correct

OpenStudy (allie2020):

ok thanks

OpenStudy (mayankdevnani):

np !